### v2

# Driving with Prior Maps: Unified Vector Prior Encoding for Autonomous Vehicle Mapping

Abstract
High-Definition Maps (HD maps) are essential for the precise navigation and decision-making of autonomous vehicles, yet their creation and upkeep present significant cost and timeliness challenges. The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather. This paper proposes the PriorDrive framework to addresses these limitations by harnessing the power of prior maps, significantly enhancing the robustness and accuracy of online HD map construction. Our approach integrates a variety of prior maps, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data. To effectively encode this prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements. At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs a dual encoding mechanism to process vector data. The intra-vector encoder captures fine-grained local features, while the inter-vector encoder integrates global context. Furthermore, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data, thereby improving the encoder’s generalizability and performance. Through extensive testing on the nuScenes dataset, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities. The integration of prior maps through the PriorDrive framework offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
Introduction
High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries (Elghazaly et al. 2023; Gao et al. 2020). Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive. These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments. In response, there is increasing interest in online HD map construction (Qiao et al. 2023; Ding et al. 2023; Zhang et al. 2024a), where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds. Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions or inclement weather conditions.
In this paper, we introduce PriorDrive, a novel framework designed to overcome these challenges by leveraging prior maps. Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure LABEL:fig:priors. SD maps provide essential long-distance centerline skeletons, which are crucial for maintaining road-level connectivity. Existing HD maps (HD map-EX), despite their high accuracy, often fail to reflect current road conditions due to infrequent updates. Nevertheless, these maps offer valuable prior information for HD map construction. Additionally, historical prediction maps, which are often overlooked, can provide insights into current road conditions by reflecting the latest observations from previous perceptions. As vehicles traverse the same locations multiple times under varying conditions, the accumulated information can be used to enhance the quality of the current local map through iterative refinement.
Recent studies (Xiong et al. 2023; Jiang et al. 2024; Sun et al. 2023) have attempted to use prior maps to supplement perception data, but they face several limitations. For instance, NMP (Xiong et al. 2023) relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments. P-MapNet (Jiang et al. 2024) incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively. MapEX (Sun et al. 2023), on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps. These actual outdated maps, due to their low update frequency, struggle to reflect real-time changes in road structures, making them unreliable for current navigation needs.
Given that various maps are typically stored as vectors due to their low redundancy and universal applicability, it’s crucial to develop a robust vector encoding model. However, designing such a model is challenging due to the variable length, sparsity, and discreteness of vector data. To address this and enhance current perception results by integrating diverse vector prior maps into various mapping models, we propose a unified vector encoder (UVE). The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow the model to fully learn the prior distribution of vector maps. Our experiments on the nuScenes dataset demonstrate that our PriorDrive can be seamlessly applied to various online mapping models, resulting in significant performance improvements, including a +2.2 mIoU for HDMapNet and a +4.2 mAP for MapTRv2.
In summary, our contributions are as follows:
  • We introduce a unified vector encoder (UVE) that effectively encodes various vector data through a dual encoding mechanism: the Intra-Vector Encoder captures fine local features, while the Inter-Vector Encoder fuses global context information.
  • We propose a pre-training paradigm for vector data that learns the prior distribution of vector data by adding Gaussian noise or mask at the segment and point levels and reconstructing the entire vector map.
  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with vector prior maps to addresses the limitations of single-perception. Our comprehensive evaluation on nuScenes dataset demonstrates that our PriorDrive significantly enhances the performance of online mapping models.
Related Work
Online Vectorized HD Map Construction
Recent advances in HD map construction have shifted from traditional offline methods based on SLAM techniques (Shan and Englot 2018; Shan et al. 2020) to more dynamic approaches that utilize vehicle-mounted sensors for online HD map generation (Xie, Pang, and Wang 2023; Chen, Deng, and Furukawa 2024; Liu et al. 2024a). This shift not only reduces costs but also ensures that maps reflect up-to-date road conditions. Some approaches (Li et al. 2021, 2022) treat HD map construction as a segmentation task, predicting pixel-level rasterized maps that require complex post-processing for vectorization. To address this limitation, recent work (Hu et al. 2024; Zhang et al. 2024b; Chen et al. 2024) has focused on developing end-to-end frameworks for vectorized map learning.
VectorMapNet (Liu et al. 2022) introduced a coarse-to-fine, two-stage network leveraging keypoint representations. MapTR (Liao et al. 2022) advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like (Carion et al. 2020) one-stage network. HIMap (Zhou et al. 2024) proposed a hybrid representation learning pipeline to capture both point-level and element-level information. MapQR (Liu et al. 2024c) introduced a scatter-and-gather query mechanism for precise detection of content and location information, while MGMap (Liu et al. 2024b) used mask-guided features to refine lane shape representations with enhanced feature detail. However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion or inclement weather conditions.
Online Mapping Based on Prior Maps
Recent research has explored the integration of prior maps to enhance the performance of online mapping models. P-MapNet (Jiang et al. 2024) encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps. NMP (Xiong et al. 2023) introduces a global neural map prior that self-updates, thereby improving the performance of local map inference. MapEX (Sun et al. 2023) categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm. Despite these advances, many studies overlook the potential of online historical prediction maps as a source of prior information. Our work addresses this gap by leveraging various existing prior maps in their native vector form to enhance current perception data.
Pretrained Methods Based on Mask Modeling
In the fields of Natural Language Processing and Computer Vision, masked pre-training has proven to be an effective strategy for self-supervised representation learning. In NLP, models like BERT (Devlin et al. 2019; Liu et al. 2019; Lan et al. 2019) use masked language modeling to predict randomly masked tokens within a bidirectional text context. Similarly, methods such as MAE (He et al. 2021; Li et al. 2023; Liu et al. 2023) mask random patches of input images and reconstruct them based on the remaining unmasked patches. Diverging from these approaches, which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data. Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized map data.
Proposed Method: PriorDrive
Figure 1: Our Integration Method. For query-based models, we enable interaction between the prior features and queries at both the instance-level and point-level, ensuring a more nuanced integration. Hierarchical query comes from MapTRv2 (Liao et al. 2023). For the other models, we integrate the prior features directly with the BEV features.
Problem Formulation
Formulation of Online Mapping.
The goal of online mapping is to construct a local HD map using on-board sensor observations, such as the surrounding camera images. The online mapping process typically involves two main modules: the Bird’s-Eye View (BEV) encoder and the map decoder. The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features. The map decoder usually comprises a Fully Convolutional Network (FCN) (Shelhamer, Long, and Darrell 2015) or a query-based module similar to DETR (Carion et al. 2020), which outputs either rasterized perception results or vectorized map elements.
Let the set of images be denoted as $\{I_{1},\ldots,I_{k}\}$, and the BEV encoder be represented as ${E}_{bev}$. The extraction of BEV features ${f}_{bev}$ can then be expressed as:
$f_{bev}={E}_{bev}(\{I_{1},\ldots,I_{K}\}).$ (1)
Similarly, let the map decoder module be denoted as $D_{map}$. The final process of predicting map elements can be expressed as:
$\mathcal{P}={D}_{map}(f_{bev},~{}opt({Q})),$ (2)
where $\mathcal{P}$ represents the predicted map elements, $Q$ denotes the learnable query, and $opt$ indicates that this learnable query is optional, which is required in query-based models but not in other models.
Extraction of Prior Map Features.
Our proposed UVE serves as a unified vector encoder that can directly encode various vector data information, such as the position ($x$, $y$), direction ($v_{x}$, $v_{y}$), and type of points ($c$). Using $p=\left[x,y,v_{x},v_{y},c\right]$ to represent the point, a vector $v$ typically consists of multiple ordered sets of points, $v=\left[p_{1},p_{2},\ldots,p_{n}\right]$, where $n$ varies. The map prior $M_{prior}$ consists of multiple vectors $M_{prior}=\{v_{1},v_{2},\ldots,v_{m}\}$, with $m$ being variable. Using $E_{uve}$ to represent the UVE model, the extraction process of the features of the prior map, denoted as $f_{prior}$, can be represented as:
$f_{prior}=E_{uve}(M_{prior}).$ (3)
Online Mapping Based on Prior Maps.
We proposed four methods in total to integrate prior map features into online mapping models. For the methods without a learnable query $Q$, such as HDMapNet (Li et al. 2021), we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{BEV}$. Then connect it with the BEV feature $f_{BEV}$ and aligned by $Conv$, as shown in Figure 1 left.
$f_{bev}^{*}=Conv(cat[f_{bev},~{}Deconv(reshape(f_{prior}))]),$ (4)
here, $f_{bev}^{*}$ represents the BEV feature after incorporating the prior map information, which can be input into the map decoder $D_{map}$ to obtain the predicted results of map elements.
For methods based on learnable query $Q$, such as the vectorized model MapTR series (Liao et al. 2023), we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 1 right. Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively. The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$. So we interact prior features and queries at the instance level and click, respectively. The HPQuery of addition, replacement, and concatenation operations can be expressed as:
$q_{ij}^{*}=q_{i}^{ins}.add(f_{i}^{ins})+q_{j}^{pt}.add(f_{j}^{pt}),$ (5)
$q_{ij}^{*}=q_{i}^{ins}.replace(f_{i}^{ins})+q_{j}^{pt}.replace(f_{j}^{pt}),$ (6)
$q_{ij}^{*}=concat[q_{i}^{ins},~{}f_{i}^{ins}]+concat[q_{j}^{pt},~{}f_{j}^{pt}].$ (7)
We will detail the effectiveness of these feature fusion methods in the experimental section.
Figure 2: Structure of UVE. Hybrid prior embedding enhances vector representation. Intra-instance and inter-instance attention mechanisms are employed to refine local features and capture global context effectively.
Architecture of UVE
Inspired by BERT (Devlin et al. 2019) on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 2). And we provide a pre-training strategy to improve the encoding and noise reduction ability of UVE.
Hybrid Prior Embedding.
As shown in Figure 2, for the given prior map $M_{prior}$, we first construct the input for UVE, termed as Hybrid Prior Embedding (HPE). We use the method from (Tancik et al. 2020) for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$. We concatenate these two embeddings as the point-level embedding. To obtain the feature representation of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding. Additionally, we introduce learnable instance embedding and learnable type embedding to distinguish vector instances. To ensure the order of points, we also introduce 2D learnable position embedding. Finally, these embeddings are aggregated to form the HPE, which then serves as the input embedding for UVE.
Intra and Inter-Vector Encoder.
Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging for the model, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector. As a result, the UVE encoder utilizes $M$-layer intra-vector attention encoders and $N$-layer inter-vector attention encoders, using an attention masking mechanism to facilitate feature interactions within and between different vector instances. Through this dual encoding mechanism, UVE can deeply understand and represent vector data, providing strong support for key tasks such as online HD maps construction.
Figure 3: Pre-trained Pipeline. The noise & mask generator creates segment-level and point-level noise or mask, and then reconstructs the entire map via UVE and MLP.
Pre-training UVE: Position Modeling
Due to the limited inference ability of online mapping model, there are errors in historical prediction maps. So we used position modeling to pre-train UVE to improve its encoding and noise reduction capabilities. specifically, there are two methods: random noise and masking (see Figure 3).
Noise & Mask Generator.
The noise is primarily categorized into segment-level and point-level. In point-level noise, random noise is added to $5\%$ of the vector points across the entire map. For segment-level noise, random noise is added to all vector points within a sub-vector segment, after randomly selecting $10\%$ of the map elements. Using $M_{org}$ to represent the original map elements (with the same data structure as $M_{prior}$), the process of adding Gaussian noise can be represented by the following formula:
$M_{org}^{*}=M_{org}\emph{[random\_index]}+\epsilon,$ (8)
where $\epsilon\sim\mathcal{N}(0,1)$ is a random noise, which is only added to the horizontal and vertical coordinates of each point.
Similar to adding noise to the points, the selection of points for adding the mask is also divided into point-level and segment-level. After selecting the points to add the mask, the coordinates of these points will be masked as $-1$.
[…]
Loss Function.
We feed the vector map to UVE encoding after passing through the noise mask generator, then decode the coordinates of all vector points using MLP, and calculate mean euclide distance from real coordinates as supervision:
$\mathcal{L}=RMSE(P,~{}mlp(E_{uve}(M_{org}^{*}))),$ (10)
where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in the $M_{org}$.
Experiments
Experimental Settings
Benchmark.
We mainly validate our method on the popular nuScenes (Caesar et al. 2019) dataset following the standard setting of previous methods (Li et al. 2021; Liu et al. 2022; Liao et al. 2022). NuScenes is a large-scale autonomous driving benchmark, that includes multiple traversals with precise localization and annotated HD maps labels. It contains a total of 1,000 scenes, each with an approximate duration of 20 seconds. The data was collected utilizing a 32-beam LiDAR operating at 20Hz and six cameras offering a 360-degree field of view at 12 Hz. Annotations for keyframes are furnished at a rate of 2Hz, with the training and validation sets containing 28,130 frames and 6,019 frames, respectively.
Evaluation Protocol.
For the vectorized results, we follow the standard metric used in previous works (Li et al. 2021; Liu et al. 2022; Liao et al. 2022). The perception range along the direction of vehicle travel are $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis. And we adopt average precision (AP) evaluated the quality of map construction. Chamfer distance $D_{Chamfer}$ is utilized to determine whether the prediction and ground truth (GT) are matched or not. We compute the $\mathrm{AP}_{\tau}$ under several $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:
$\mathrm{AP}=\frac{1}{|\mathrm{T}|}\sum_{\tau\in\mathrm{T}}\mathrm{AP}_{\tau}.$ (11)
Following the previous methods (Li et al. 2021; Liu et al. 2022; Liao et al. 2022), we have selected three kinds of map elements are chosen for fair evaluation: lane divider, pedestrian crossing and road boundary. Furthermore, to evaluate the effectiveness of our method across different evaluation metrics and for the rasterized results, we follow the previous works (Li et al. 2021; Xiong et al. 2023; Jiang et al. 2024) and employ intersection-over-union (IoU) as the metrics for segmentation results.
Implementation Details.
We primarily perform our experiments using the latest and influential vectorized model MapTRv2 (Liao et al. 2023). In order to validate the broad applicability of our method, we also use HDMapNet (Li et al. 2021) as baseline. For fair comparisons, experiments were conducted on all baseline models using their original configs and parameters without any changes. All experiments were trained with 6 NVIDIA RTX A6000. For pre-training UVE, we trained the 24-epoch on nuScenes vecror data. For historical prediction maps, we run each base model to get their respective online local maps.
Method Modality Backbone Epoch Metric ped. div. bou. mean $\mathbf{\tau_{0.5m}}$
HDMapNet C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
+ SD map Prior C Effi-B0 30 IoU 20.8 41.7 41.8 34.8 -
+ Online Local Map Prior C Effi-B0 30 IoU 22.5 41.3 41.7 35.1 -
+ HD map-EX* Prior C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
MapTRv2 C R50 24 AP 59.8 62.4 62.4 61.5 38.6
+ SD map Prior C R50 24 AP 61.7 65.5 66.4 64.5 42.7
+ Online Local Map Prior C R50 24 AP 63.6 67.3 66.2 65.7 44.7
+ HD map-EX* Prior C R50 24 AP 61.5 65.6 68.3 65.1 44.3
Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range. “C” denotes camera input. The best results using the same backbone are highlighted in bold. The SD map is sourced from OpenStreetMap (Haklay and Weber 2008). The online local map refers to the historical prediction results. As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX (Sun et al. 2023) to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.
M N Dim IoU
Intra Inter ped. div. bou. mean
4 4 64 20.7 41.3 41.1 34.4
6 2 64 20.5 41.1 40.9 34.2
2 6 64 20.7 41.0 41.3 34.3
2 2 128 19.5 41.4 41.2 34.0
2 2 64 20.8 41.7 41.8 34.8
1 1 32 19.8 41.1 40.9 33.9
Table 2: Ablation experiments on UVE structure within HDMapNet and SD map prior.
Method Dist Err. AP
Type Seg Pt ped. div. bou. mean
Without Pretrain - 61.3 63.4 64.4 63.0
Original - - 0.09m 61.2 64.1 64.5 63.3
Noise 3m 20% - 0.46m 61.4 63.0 63.8 62.7
Noise 1m 10% - 0.15m 61.6 63.8 64.2 63.2
Noise 1m 10% 5% 0.18m 61.6 64.3 64.8 63.6
Mask 10% 5% 0.43m 61.8 64.0 64.3 63.4
Table 3: Ablation experiments of the pre-training strategies on MapTRv2. Here, we fixed prior_num=2, search_range=5m and integration=replace. “Seg” and “Pt” represent segment-level and point-level, respectively.
Main Results
To evaluate the effectiveness of our method across different model architectures, evaluation metrics, and vector prior maps, we integrated the UVE-encoded prior maps into both HDMapNet and the vectorized model MapTRv2 (see Table 1). The prior information provided by SD maps, which primarily consists of a centerline skeleton, resulted in the smallest improvement. Specifically, MapTRv2 and HDMapNet showed enhancements of 3.0 mAP and 1.9 mIoU, respectively, when using SD map priors. And for query-based model, our method has a greater improvement because our HPQuery makes better use of prior information. The incorporation of online local maps, which offer up-to-date information, led to more significant improvements. This is reflected in the enhancements of 4.2 mAP and 2.2 mIoU for MapTRv2 and HDMapNet, respectively. These results indicate that more accurate and current local maps contribute to better performance by providing more comprehensive and precise information. Figure 4 illustrates how online local maps effectively restore obscured lane dividers, pedestrian crossings, and road boundaries. Although offline HD maps are highly accurate, their lack of complete prior information resulted in notable improvements of 3.6 mAP and 3.2 mIoU for MapTRv2 and HDMapNet, respectively. Additionally, we evaluated the results using a stricter threshold of 0.5m. Under this criterion, our method with SD map, online local map, and outdated HD map (HD map-EX*) priors achieved improvements of 4.1 mAP, 6.1 mAP, and 5.7 mAP, respectively. These results underscore that our method provides more substantial improvements under stricter evaluation conditions. Overall, our method demonstrated consistent performance improvements across all baseline models with various prior maps, highlighting its generalizability and potential applicability to other mapping frameworks.
Ablation Study
In this section, we do extensive research on UVE structure, pre-training, number and search range of prior maps, and integration methods. we conduct experiments using online local map priors on MapTRv2 unless otherwise specified.
Structure of UVE.
Table 2 presents the results of ablation experiments on the UVE structure using the SD prior map within HDMapNet. We observed that increasing the number of layers or expanding the feature dimension of UVE yielded only marginal improvements, averaging around 1.3 mIoU. This limited gain may be attributed to the model’s excess capacity, leading to overfitting. Conversely, a minimal configuration of UVE, with only one layer each of intra-vector and inter-vector attention and a dimension of 32, resulted in a modest improvement of just 1 mIoU. However, when UVE was configured with two layers of intra-vector and inter-vector attention and a feature dimension of 64, the model’s performance reached saturation.
The Effect of Pre-training.
To effectively capture the prior information of vector maps, we pre-trained UVE using different strategies, as shown in Table 3. When we fed the original vector map directly to UVE without any noise, UVE converged with a minimal error of 0.09 meters. However, the simplicity of the task might have limited the model’s ability to fully learn the prior distribution, resulting in only a 0.3 mAP improvement. We added segment-level Gaussian noise with a 3-meter standard deviation to 20% of the map elements randomly. The UVE’s average Euclidean distance error in reconstructing the entire map was 0.46 meters. However, this approach led to a slight decrease in mAP, likely because the UVE had not fully converged. Next, we applied 1-meter segment-level noise to 10% of the map elements randomly, resulting in a reduced error of 0.15 meters for UVE and a 0.2 mAP improvement compared to the model without pre-training. Building on this, we introduced 1-meter point-level noise to 5% of the vector points. Although this increased UVE’s error to 0.18 meters, it yielded a more substantial mAP improvement of 0.6. This demonstrates that pre-training strategies incorporating both segment-level and point-level noise are more effective at learning the prior information of vector maps. We also experimented with masking instead of noise, which increased UVE’s error to 0.43 meters but still resulted in a 0.4 mAP improvement. This suggests that while challenging tasks based on masking can pre-train UVE effectively, their impact may be constrained by the amount of pre-training data.
Impact of the Number of Prior Maps.
We conducted an ablation study to assess the effect of varying the number of prior maps on model performance (see Table 4). The results indicate that using a single prior map led to a 0.9 mAP improvement, demonstrating its effectiveness in enhancing current perception. Incorporating two prior maps resulted in a more substantial improvement of 1.5 mAP, reflecting the value of richer prior information. However, using multiple prior maps only yielded a 1.3 mAP improvement. This could be attributed to the integration method, where replacing too many original queries with prior features may have diminished the performance of the original queries.
Effect of Different Search Ranges for Prior Maps.
We performed ablation experiments to explore the impact of search ranges for prior maps in Table 5. The results indicate that increasing the search range leads to a slight decrease in improvement. This suggests that while a larger search range allows the prior map to cover a wider area, it also introduces greater positional error. On the other hand, a more precise, narrower search range may not always provide multiple prior maps for most locations, limiting the benefits.
Different Methods for Integrating Prior Maps.
To evaluate the versatility of our approach, we tested various integration methods in Table 6. Compared to MapTRv2, the HPQuery methods of replacement and addition resulted in mAP improvements of 2.1 and 1.6, respectively, but these methods may impact the performance of the original query. In contrast, the concatenation method achieved a more substantial improvement of 4.2 mAP. This approach effectively preserves the original query’s integrity while also leveraging the instance-level and point-level prior features learned by UVE, making it the most effective integration method.
Inference Speed, Memory, Model Size, and Detailed Runtime.
Table 7 compares our method with MapTRv2 in the above aspects. The results show that our method maintains similar FPS, GPU memory usage, and parameter count compared to MapTRv2, while achieving a 4.2 mAP improvement. Additionally, we provide the detailed runtime of each component. Our proposed UVE accounts for only 15.3% of the total pipeline runtime, with the majority of inference time spent on the 2D Backbone (50.2%), highlighting the efficiency and practical value of our approach across various application scenarios.
Fixed Prior AP
Parameters Num ped. div. bou. mean
- None 59.8 62.4 62.4 61.5
search_range=5m One 61.9 62.1 62.4 62.4
pretrain=false Two 61.3 63.4 64.4 63.0
integration=replace Multiple 61.2 63.0 64.2 62.8
Table 4: Ablation experiments on the number of online local prior maps on MapTRv2.
Fixed Search AP
Parameters Range ped. div. bou. mean
- None 59.8 62.4 62.4 61.5
prior_num=2 5 m 61.3 63.4 64.4 63.0
pretrain=false 10 m 60.8 63.7 64.0 62.8
integration=replace 15 m 61.5 62.7 64.0 62.7
Table 5: Ablation experiments on the search range of online local prior maps on MapTRv2.
Fixed Integrated AP
Parameters Method ped. div. bou. mean
- None 59.8 62.4 62.4 61.5
prior_num=2 Replace 61.6 64.3 64.8 63.6
pretrain=true Add 62.5 63.3 63.6 63.1
search_range=5m Concat 63.6 67.3 66.2 65.7
Table 6: Ablation experiments on the integrated methods of online local prior maps on MapTRv2. Note that we proposed other integrated methods after pre-training UVE, so we use pre-trained UVE here.
- MapTRv2 Ours
FPS 10.3 9.9
GPU mem. (MB) 2624 2650
Params (MB) 40.3 43.4
mAP 61.5 65.7
Component Runtime (ms) Proportion
2D Backbone 47.6 50.2%
PV to BEV Module 12.9 13.6%
UVE 14.6 15.3%
Map Decoder 19.8 20.9%
Table 7: Detailed comparison of our method with MapTRv2. FPS and detailed runtime of each component are measured on one A6000 GPU with batch size as 1.
Figure 4: Qualitative results with and w/o online local prior. Prior maps can help restore obscured map elements and make predictions more complete and accurate.
Conclusion
In this paper, we introduced PriorDrive, a novel framework that effectively leverages various types of prior maps to enhance the accuracy and robustness of online HD map construction for autonomous vehicles. Central to our approach is the UVE, which we designed to encode diverse vector data efficiently. Through comprehensive experimentation, we demonstrated that UVE, combined with our proposed pre-training strategies, significantly improves the performance of state-of-the-art mapping models. Our approach not only addresses the challenges associated with real-time HD map construction in dynamic and complex environments but also offers a scalable solution that continually improves map accuracy over time. The iterative use of historical prediction maps as priors leads to progressively refined map outputs, thus ensuring that the most up-to-date road information is available for autonomous navigation.